OpenStudy (anonymous):

Solve for x: |x + 9| = 17 A. {8, 26} B. {8, -26} C. {-8, 26} D. {-8, -26}

OpenStudy (anonymous):

First break |x + 9| up as a piece-wise defined function: 1.) |x + 9| equals x + 9 when x + 9 ≥ 0. 2.) |x + 9| equals -(x + 9) when x + 9 < 0. Now that you've set a range, you can get rid of the absolute value and solve for x in both cases: 1.) x + 9 = 17 → x = 17 - 9 → x = 8 2.) -(x + 9) = 17 → -x - 9 = 17 → -x = 26 → x = -26 Your solutions are 8 and -26, which is B.

OpenStudy (anonymous):

x+9=17 , x+9=-17 x=8 , x=-26 the answer is B. {8, -26}